If a pension fund is looking at a certain European put option with four months to expiration on a non-dividend paying stock. If the current stock price is $19.29 with a strike price of $21.25 and the risk-free interest rate is 1.75% per annum, what is a lower bound for the price of this option?
Answer #1

Lower bound = Stock Price - Strike Price*e-rt =19.29-21.25*EXP(-1.75%*4/12)=-1.84